\begin{thebibliography}{00}
	
\bibitem{Carmo-Eidt-Noor}  J. R. Carmo, Ben H. Eidt and S. W. Noor. \textit{Minimal invariant subspaces for an affine composition operator}. Complex Anal. Oper. Theory 18, 59 (2024). https://doi.org/10.1007/s11785-024-01501-9.
   \bibitem{Beurlingtype} S. Bose, P. Muthukumar and J. Sarkar. \textit{Beurling type invariant subspaces of composition operators}. J. Operator Theory, 86(2) (2021) 137-156.

    \bibitem{Modeltype} S. Bose, P. Muthukumar and J. Sarkar. \textit{Model spaces invariant under composition operators}. Canad. Math. Bull. Vol. 66 (1), 2023, 204–217.

	\bibitem{shapirocyclic} P. Bourdon and J. H. Shapiro, \textit{Cyclic phenomena for composition operators}. Mem. Amer. Math. Soc., 596 (1997).

   % \bibitem{Posinormal Bourdon} P. Bourdon and D. Thompson. \textit{Posinormal composition operators on $H^2$}. Journal of Mathematical Analysis and Applications. 518.2 (2023): 126709.

	% \bibitem{caradus} S. R. Caradus. \textit{Universal operators and invariant subspaces}. Proc. Amer. Math. Soc. 23 (1969): 526-527. 
	
	\bibitem{noor} J. R. Carmo and S. W. Noor. \textit{Universal composition operators}. J. Operator Theory, 87 (2022) 137-156.
	
	% \bibitem{modernisp} I. Chalendar and J. R. Partington. \textit{Modern approaches to the Invariant Subspace Problem}. 188, Cambridge University Press (2011).
	
	%\bibitem{Chkliar} V. Chkliar. \textit{Eigenfunctions of the hyperbolic composition operator}. Integral Equations Operator Theory, 29 (1997): 364-367.

    % \bibitem{CowenGallardo1} C. C. Cowen and Eva A. Gallardo Gutiérrez. \textit{Consequences of Universality Among Toeplitz Operators}. J. Math. Anal. Appl., 432.1 (2015): 484–503.
   

	%\bibitem{Cowen-Maccluer} C. C. Cowen and B. MacCluer. \textit{Composition Operator on Spaces of Analytic Functions}. Routledge (2019).
 
 %Studies in Advanced Mathematics. CRC Press, Boca Raton, 1995.

    \bibitem{Cowen-Wahl} C. C. Cowen and R. G. Wahl, \textit{Shift-invariant subspaces invariant for composition operators on the Hardy-Hilbert space}. Proc. Amer. Math. Soc. 142.12 (2014): 4143-4154.

	

    \bibitem{fricain1} E. Fricain and Mashreghi, J. \textit{The theory of $H(b)$ spaces}. Vol 1. Cambridge University Press (2016). 
 
	
	%\bibitem{Gallardo-Gorkin} E. A. Gallardo-Guti\'{e}rrez, P. Gorkin, \textit{Minimal invariant subspaces for composition operators}, J. Math. Pures Appl. 95 (2011) 245-259.

    \bibitem{Cesaro 2} E. A Gallardo-Guti\'{e}rrez, J. R. Partington and W. T. Ross. \textit{Invariant subspaces of the Cesaro operator}. arXiv preprint arXiv:2307.06923 (2023).  

    \bibitem{Cesaro 1} E. A Gallardo-Guti\'{e}rrez and J. R. Partington. \textit{Insights on the Ces{\`a}ro operator: shift semigroups and invariant subspaces}. Journal d'Analyse Math., 152 (2024), 595-614.

	%\bibitem{Hurst} P. R. Hurst. \textit{Relating composition operators on different weighted Hardy spaces}. Arch. Math. 68 (1997): 503-513.
	
	\bibitem{hhspaces} R. A. Mart{\'\i}nez-Avenda{\~n}o and P. Rosenthal. \textit{An introduction to operators on the Hardy-Hilbert space}. Graduate Texts in Mathematics (GTM, volume 237), Springer (2007).
 
	%\bibitem{Matache minimal} V. Matache. \textit{ On the minimal invariant subspaces of the hyperbolic composition operator}. Proc. Amer. Math. Soc. 119 (3) (1993) 837-841.
	
	% \bibitem{Matache eigenfunction} V. Matache. \textit{The eigenfunctions of a certain composition operator}. Contemp. Math. 213 (1998): 121-136.
	
	
	%\bibitem{Mortini} R. Mortini. \textit{Cyclic subspaces and eigenvectors of the hyperbolic composition operator}; Sém. Math. Luxembourg, in: Travaux Mathématiques, Fasc. VII, Centre Univ. Luxembourg, Luxembourg, (1995): 69–79.
	
	%\bibitem{Nordgren-Rosenthal-Wintrobe} E. Nordgren, P. Rosenthal, F.S. Wintrobe. \textit{Invertible composition operators on $H^p$} J. Funct. Anal. 73 (1987): 324-344.
	
	
	%\bibitem{pozzi} E. Pozzi. \textit{ Universality of weighted composition operators on $L^2([0,1])$ and Sobolev spaces}. Acta Sci. Math., 78 (2012): 609-642.
	
	\bibitem{rota} G. C. Rota. \textit{On models for linear operators}. Comm. Pure Appl. Math. 13 (1960): 469-472.

	
	% \bibitem{shapiro} J. H. Shapiro. \textit{The essential norm of a composition operator}. Ann. of Math. 125, (1987): 375-404.
	
	% \bibitem{Shapiro Text} J. H. Shapiro, \textit{Composition operators and classical function theory}. Universitext, Springer-Verlag (1993).
	
	
	% \bibitem{Wogen} W. R. Wogen. \textit{On some operators with cyclic vectors}. Indiana Univ. Math. J., 27 (1978): 163-171.

  
\end{thebibliography}

\documentclass{amsart}
\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}

\newtheorem{thm}{Theorem}
\newtheorem{cor}[thm]{Corollary}
\newtheorem{lem}[thm]{Lemma}
\newtheorem{prop}[thm]{Proposition} 
\theoremstyle{definition}

\newtheorem{exem}[thm]{Example}
\newtheorem*{theorem*}{Theorem}

\usepackage{float}
\usepackage{tikz}
\usetikzlibrary{shapes,shapes.geometric,arrows,fit,calc,positioning,automata}

\theoremstyle{remark}

\numberwithin{equation}{section}

\def\<{\langle}
\def\>{\rangle}
\begin{document}
\title[]{Beurling and Model subspaces invariant under a universal operator}
\author{Ben Hur Eidt and  S. Waleed Noor}%
\address{IMECC, Universidade Estadual de Campinas, Campinas-SP, Brazil.}
\email{$\mathrm{b264387@dac.unicamp.br}$ (Ben Hur Eidt)} 
\email{$\mathrm{waleed@unicamp.br}$ (Waleed Noor) Corresponding Author}

\begin{abstract}  In this article, we characterize the Beurling and Model subspaces of the Hardy-Hilbert space $H^2(\mathbb{D})$  invariant under the composition operator $C_{\phi_a}f=f\circ\phi_a$, where $\phi_a(z) = az + 1 - a$ for $a \in (0,1)$ is an affine self-map of the open unit disk $\mathbb{D}$. These operators have universal translates (in the sense of Rota) and have attracted attention recently due to their connection with the Invariant Subspace Problem (ISP) and the classical Cesàro operator.

\end{abstract}

{\subjclass[2010]{Primary; Secondary}}
\keywords{Composition operators, model spaces, Beurling type spaces, invariant subspace problem.}
\maketitle{}

\section{Introduction} 

The Hardy-Hilbert space of the open unit disk $\mathbb{D}$, denoted by $H^2$ is the Hilbert space of holomorphic functions $f : \mathbb{D} \to \mathbb{C}$ such that 
$$\|f\|^2 = \sup\limits_{0 < r < 1} \frac{1}{2\pi} \int\limits_{0}^{2\pi} |f(re^{i \theta} )|^2 d\theta < \infty.$$
 If $\phi:\mathbb{D} \to \mathbb{D}$ is a holomorphic self-map of $\mathbb{D}$, then $C_{\phi}f = f \circ \phi$ is the composition operator with symbol $\phi$. The Littlewood subordination Theorem ensures that $C_{\phi}$ is always a bounded linear operator on $H^2$. The study of composition operators centers around the interaction between the function-theoretic properties of the symbol $\phi$ and the operator-theoretic properties of $C_\phi$.

The \emph{Invariant Subspace Problem} (ISP) can be stated as follows: Does every bounded linear operator on complex separable Hilbert space have a non-trivial closed invariant subspace? In this article we are interested in the affine symbols $\phi_a(z) = az + 1 - a$ for $a \in (0,1)$. In \cite{noor} the authors show that $C_{\phi_a} - \lambda$ is \emph{universal} in the sense of Rota (see \cite{rota}) for some numbers $\lambda \in \mathbb{C}$ and how it can be used to reformulate the ISP:

\begin{center}
   \textit{The ISP is true if and only if every minimal invariant subspace of $C_{\phi_a}$ is one-dimensional.} 
\end{center}

All invariant subspaces are assumed to be closed, and \emph{minimality} here imples that it does not contain another proper invariant subspace. Results about minimal \emph{cyclic} invariant subspaces were obtained recently in \cite{Carmo-Eidt-Noor}. The composition operators $C_{\phi_a}$ also appeared recently in \cite{Cesaro 2} and \cite{Cesaro 1}  where the authors considered the holomorphic flow given by $\varphi_t(z) = e^{-t}z + 1 - e^{-t}$ for $t \geq 0$. The focus here is on the classical Césaro operator defined on $H^2$ by
$$(\mathcal{C}f)(z) = \sum\limits_{n = 0}^{\infty} \left( \frac{1}{n + 1}\sum\limits_{k = 0}^na_k \right) z^n$$
where $(a_n)_{n\in\mathbb{N}}$ are the Taylor coefficients of $f\in H^2$. A complete description of the invariant subspaces of the Cesàro operator remains an open problem in concrete operator theory. A one-to-one correspondence between the invariant subspaces of $\mathcal{C}$ and the \emph{common} invariant subspaces of the family $\Phi = (\varphi_t)_{t \geq 0}$ is established in \cite[Theorem $2.1$]{Cesaro 1}. 

\begin{thm} \label{GP1}
A closed subspace $M$ of $H^2$ is invariant under the Cesàro operator $\mathcal{C}$ if and only if its orthogonal complement $M^{\perp}$ is invariant under the semigroup $\Phi$.
\end{thm}

It is well-known that for any $f \in H^2$ the radial limit 
$$f^*(e^{i \theta}) := \lim\limits_{r \to 1^{-}} f(re^{i \theta})$$
exists almost everywhere with respect to the normalized Lebesgue measure on $\mathbb{T}$ and the correspondence $f \to f^*$ gives us an isometric isomorphism between $H^2$ and a subspace of $L^2(\mathbb{T})$. See \cite{hhspaces} for more details. A function $\Theta \in H^2$ such that $|\Theta^*(e^{i \theta})| = 1$ almost everywhere in $\mathbb{T}$ is called an \textit{inner function}. The celebrated Beurling's Theorem states that a subspace $M \subseteq H^2$ is shift-invariant (i.e $zM\subset M$) if and only if $M = \Theta H^2$ for some inner function $\Theta$. These subspaces are called the \emph{Beurling subspaces}. Their orthogonal complements in $H^2$, i.e, the subspaces $(\Theta H^2)^{\perp}$ are called \textit{model spaces}. Recently, the invariant Beurling and model subspaces for composition operators were investigated by Bose, Muthukumar and Sarkar (see \cite{Beurlingtype} and \cite{Modeltype}). 

The main goal of this article is to characterize the model and Beurling subspaces of $H^2$ that are invariant under $C_{\phi_a}$ for some $a \in (0,1)$. The paper is organized as follows. In Section 2 we present some preliminaries.
In Section 3 we characterize all the model spaces that are invariant under $C_{\phi_a}$. The main result of this section states that a model space $(\Theta H^2)^{\perp}$ is invariant under $C_{\phi_a}$ if and only if $\Theta(z) = z^n$ for some $n \in \mathbb{N}_{0}$ (see Theorem \ref{mp}). In Section $4$, we consider the Beurling subspaces $\Theta H^2$ that are invariant under $C_{\phi_a}$. Here we obtain the following dichotomy: $\Theta $ is either the atomic singular inner function or $\Theta$ has infinitely many zeros accumulating at $1$ (see Theorem \ref{dic}).  As an immediate consequence of these results, Theorem \ref{GP1} provides characterizations of the invariant Beurling and model subspaces  for the Cesàro operator (see Corollaries \ref{Cesaro Beurling} and \ref{Cesaro Model}). The latter result was obtained recently by Gallardo-Gutiérrez, Partington and Ross \cite[Theorem 7.7]{Cesaro 2}.

\section{Preliminaries}

\subsection{Spectra of Schür functions}Denote by $H^{\infty}$  the space of all bounded analytic functions on $\mathbb{D}$ and denoted by $\|f\|_{\infty}$ the sup-norm of $f\in H^\infty$. The closed unit ball of $H^{\infty}$ is denoted by $\mathcal{S}:=\mathcal{S}(\mathbb{D})=\{f \in H^{\infty} \,\ | \,\ \|f\|_{\infty} \leq 1 \}$ and is called the \textit{Schür Class}. If $f \in \mathcal{S}$ we say that a point $z \in \overline
{\mathbb{D}}$ is a \textit{regular} point for $f$ if $z \in \mathbb{D}$ and $f(z) \neq 0$ or if $z \in \mathbb{T}$ and $f$ admits an analytic continuation across a neighborhood $V$ of $z$ with $|f| = 1$ in $V \cap \mathbb{T}$. The set of all regular points of $f$ is denoted by $\rho(f)$ and the set $\overline{\mathbb{D}} - \rho(f)$ is called the \textit{spectra} of $f$ and denoted by $\sigma(f)$. If $\Theta$ is an inner function and $\Theta$ admits an analytic continuation at some neighborhood $V$ of $z \in \mathbb{T}$ then the continuity of $\Theta$ in $V \cap S^1$ and the fact that $|\Theta^*(e^{i \theta})| = 1$ almost everywhere implies that $|\Theta| = 1$ in $V \cap S^1$. Every inner function $\Theta \in H^2$ can be written as $\Theta = \lambda BS_{\mu}$ where $\lambda$ is a unimodular constant, $B$ is the Blaschke product formed by the zeros of $\Theta$ and $S_{\mu}$ is a singular inner function associated to a measure $\mu$. More specifically:
$$S_{\mu}(z) = \exp \left( - \int\limits_{S^1} \frac{\xi+ z}{\xi - z} d\mu(\xi) \right).$$
where $\mu$ is a finite positive Borel measure on $\mathbb{T}$ which is singular with respect to the Lebesgue measure. For any $\xi \in \mathbb{T}$ and $\epsilon > 0$ consider the open arc of length $2\epsilon$ with center $\xi$ given by $A(\xi, \epsilon) := \{ \xi e^{it} \,\ | -\epsilon < t < \epsilon \}$. The \textit{support} of the measure $\mu$ is denoted $\mathrm{supp}(\mu) := \{\xi \in \mathbb{T} \,\ | \,\ \mu(A(\xi,\epsilon)) > 0 \,\ \,\ \forall \epsilon > 0\}.$ We refer the reader to the texts \cite{fricain1} and \cite{hhspaces} for more details. We shall need the following characterization for the spectra of an inner function.

\begin{prop}\cite[Theorem 5.4]{fricain1} \label{spectrainner}
   Let $\Theta$ be a non-constant inner function and let $\Theta(z) = \lambda B S_{\mu}$ be the canonical factorization of $\Theta$ where $\lambda\in\mathbb{T}$, $B$ is the Blaschke factor and $S_{\mu}$ is the singular inner part associated with the measure $\mu$.  Then
   $$\sigma(\Theta) = \left\{z \in \overline{\mathbb{D}} \,\ | \,\ \liminf\limits_{w \to z, \,\  w \in \mathbb{D}} |\Theta(w)| = 0 \right\} = \overline{Z(\Theta)} \cup \mathrm{supp}(\mu).$$
   where $Z(\Theta)$ denotes the set of all zeros of $\Theta$ in $\mathbb{D}$.
\end{prop}

\section{Invariant Model spaces}

 In this section our goal is to characterize the $C_{\phi_a}$-invariant model spaces $(\Theta H^2)^{\perp}$ where $\Theta \in H^2$ is an inner function. The simplest examples are the polynomials of degree at most $N$ denoted by $\mathbb{C}_N[z]$ which form the model space $(e_{N+1} H^2)^{\perp}$ where $e_n(z) = z^n$ for $n\in\mathbb{N}$. Indeed, since $C_{\phi_a}^n = C_{\phi_{a^n}}$ for $a \in (0,1)$ and $n\in\mathbb{N}$, it is easy to see that $C_{\phi_a}^np\in \mathbb{C}_N[z]$ whenever $p\in\mathbb{C}_N[z]$. We shall prove that these are infact the only examples. We begin  with some lemmas.

\begin{lem}\label{lemmamodel}
    Let $a \in (0,1)$. Then for all $\theta \in (0,2\pi]$

    $$\left| \frac{ae^{i \theta}}{1 - e^{i \theta} + ae^{i \theta}} \right| \leq 1$$
    and if

    $$\left| \frac{ae^{i \theta}}{1 - e^{i \theta} + ae^{i \theta}} \right| = 1$$ for some $\theta \in (0,2\pi]$ then $e^{i \theta} = 1$.
\end{lem}

\begin{proof}
    This follows by a computation writing $e^{i\theta} = x + iy$ with $x^2 + y^2 = 1$.
\end{proof}
For each $w \in \mathbb{D}$, let $\kappa_w(z) = \frac{1}{1 - \overline{w}z}$ be the \textit{reproducing kernel} at $w$ which satisfies $\langle f, \kappa_w \rangle = f(w)$ whenever $f \in H^2$. The next lemma can be found in \cite{Carmo-Eidt-Noor}.


\begin{lem}\label{rparecyclic}
    Let $\kappa_\alpha \in H^2$ be a reproducing kernel. Then $\kappa_\alpha$ is a cyclic vector for $C_{\phi_a}$ if and only if $\alpha \neq 0$.
\end{lem}

The next two resuts are central to our main theorem.
\begin{lem}
    Let $\Theta$ be a non-constant inner function. If $(\Theta H^2)^{\perp}$ is $C_{\phi_a}$-invariant then $\Theta(z_1) \neq 0$ for all $z_1 \in \mathbb{D} - \{0\}$.
\end{lem}
\begin{proof}
    For the sake of contradiction, suppose that $\Theta(z_1) = 0$ for some $z_1 \in \mathbb{D} - \{0\}$. Thus
  $$\langle \Theta f, \kappa_{z_1} \rangle = \Theta(z_1)f(z_1) = 0 \,\ \forall f \in H^2$$
   which implies $\kappa_{z_1} \in  (\Theta H^2)^{\perp}$. By hypothesis, this is a $C_{\phi_a}$-invariant subspace, so $K_{\kappa_{z_1}} \subseteq (\Theta H^2)^{\perp}$. By Lemma \ref{rparecyclic} $\kappa_{z_1}$ is a cyclic vector (because $z_1 \neq 0$) and thus $H^2 = (\Theta H^2)^{\perp}$ which implies $\{0\} = \Theta H^2$. So $\Theta$ is constant which give us a contradiction. We conclude that $\Theta$ does not have zeros in $\mathbb{D} - \{0\}$ as desired. 
\end{proof}

\begin{prop}\label{m2}
    Let $M = (\Theta H^2)^{\perp}$ be a model space where $\Theta $ is a singular inner function. Then $M$ is not $C_{\phi_a}$-invariant.
\end{prop}

\begin{proof}

Suppose that $M = (\Theta H^2)^{\perp}$ is $C_{\phi_a}$-invariant, we will arrive at a contradiction. Consider the function $\sigma(z) = \frac{az}{1 - (1 - a)z}$ which is a holomorphic self map of $\mathbb{D}$. Note that
$${(\Theta \circ \sigma)^{*}(e^{i \theta})}  = \lim\limits_{r
   \to 1^{-}} \Theta \circ \sigma(re^{i \theta}) =  \lim\limits_{r 
   \to 1^{-}} \Theta( \sigma(re^{i \theta} )) =  \lim\limits_{r 
   \to 1^{-}} \Theta \left( \frac{are^{i \theta}}{1 - (1 - a)re^{i\theta}} \right).$$
   If $\theta \neq 2\pi$ then by Lemma \ref{lemmamodel} we have $\left( \frac{ae^{i \theta}}{1 - (1 - a)e^{i\theta}} \right) \in \mathbb{D}$. Using that $\Theta$ is continuous in $\mathbb{D}$ and $|\Theta(w)| < 1$ for all $w \in \mathbb{D}$ (because $\Theta$ is non-constant)  we obtain:

   $$|{(\Theta \circ \sigma)^{*}(e^{i \theta})}| \overset{a.e}{=} \left|\Theta \left( \frac{ae^{i \theta}}{1 - (1 - a)e^{i\theta}} \right) \right| \overset{a.e}{<} 1.$$
  By (\cite{Modeltype}, Theorem 4.3) $M$ is invariant under $C_{\phi_{a}}$ if, and only if, $\Theta H^2$ is invariant under $C_{\sigma}$. Thus there exists $g \in H^2$ such that 
\begin{align}
      \Theta \circ \sigma = C_{\sigma}(\Theta) = \Theta g \label{x1} \implies \Theta(\sigma(z)) = \Theta(z)g(z) \,\ \,\ \forall z \in \mathbb{D}.
  \end{align}
  Passing to the radial limits and considering the modulus we conclude that 
$$1 \overset{a.e}{>} |{(\Theta \circ \sigma)^{*}(e^{i \theta})}| \overset{a.e}{=}  |g^*(e^{i \theta})|.$$
    As $g \in H^2$ this implies that $|g(z)| \leq 1$ for every $z \in \mathbb{D}$ see for example (\cite{hhspaces}, pg.14, Corollary 1.1.24). But, if we consider $z = 0$ in (\ref{x1}) we conclude that 
    \begin{align*}
        \Theta(0) = \Theta (\sigma(0)) = \Theta(0) g(0) 
    \end{align*}
    and thus g(0) = 1 because $\Theta$ is inner singular, in particular, zero-free. So by the maximum module principle, $g$ is constant and $g \equiv g(0) = 1$. Looking at (\ref{x1}) again we conclude that $C_{\sigma}(\Theta) = \Theta$ and then $\Theta$ is a fixed point. But considering the equality
$$ \Theta \circ \sigma = C_{\sigma}(\Theta) = \Theta,$$
   passing to radial limits and using the estimates proved above we obtain
$$1 \overset{a.e}{>} | (\Theta \circ \sigma)^{*}(e^{i \theta}) | \overset{a.e}{=} | \Theta^*(e^{i \theta})| \overset{a.e}{=} 1 $$
   which is a contradiction. So $\Theta$ is constant and we arrived at a contradiction.
   \end{proof}

\begin{cor}
    Let $\Theta$ be an inner function. If $\Theta(z) = z^nS(z)$ for some singular inner $S$ and $n \geq 1$, then $(\Theta H^2)^{\perp}$ is not $C_{\phi_a}$-invariant.
\end{cor}

\begin{proof}
     By \cite[Theorem 4.3]{Modeltype} $(\Theta H^2)^{\perp}$ is invariant under $C_{\phi_{a}}$ if and only if $\frac{\Theta \circ \sigma}{\Theta} \in H^{\infty}$ where $\sigma(z) = \frac{az}{1 - (1 - a)z}$. Note that

    $$ \frac{\Theta \circ \sigma (z)}{\Theta(z)} =  \frac{a^n z^n  S \circ \sigma (z)}{ (1 - (1 - a)z)^n z^n S(z)} = \frac{a^n S \circ \sigma (z)}{ (1 - (1 - a)z)^n S(z)}.$$
    If this function belongs to $H^{\infty}$ then $\frac{ S \circ \sigma}{S} \in H^{\infty}$ and this implies that $(SH^2)^{\perp}$ is $C_{\phi_a}$ invariant using again (\cite{Modeltype}, Theorem 4.3). This contradicts Theorem \ref{m2}.
\end{proof}

Putting these results together, we arrive at the main theorem of this section.

\begin{thm}\label{mp}
    The only model spaces that are invariant under $C_{\phi_a}$ for any $a\in(0,1)$ are of the form $(e_nH^2)^{\perp}$ for some $n \in \mathbb{N}$.
\end{thm}
Since this characterization is independent of $a\in(0,1)$, Theorem \ref{GP1} immediately gives the following.
\begin{cor}\label{Cesaro Beurling}
    A Beurling subspace $\Theta H^2$ is invariant under the Cesàro operator if and only if $\Theta(z)=z^n$ for some $n \in \mathbb{N}$.
\end{cor}

In what follows we will explore more details about the spaces $(z^nSH^2)^{\perp}$ and their relation with $C_{\phi_a}$. The reproducing kernels of the model space $(\Theta H^2)^{\perp}$ are given by the functions
$$\kappa_{\lambda}^{\Theta}(z) = \frac{1 - \overline{\Theta(\lambda)} \Theta(z)}{1 - \overline{\lambda}z} \,\ \,\ \,\ $$
where $\lambda\in\mathbb{D}$ (see \cite[Corollary $14.12$]{fricain1}).

\begin{prop}
    If $\Theta$ is inner and non constant, the space $z^n(\Theta H^2)^{\perp}$ is not $C_{\phi_a}$-invariant.
\end{prop}
\begin{proof}
    Suppose that this space is invariant and let $z^ng \in z^n(\Theta H^2)^{\perp}$ where $g \in (\Theta H^2)^{\perp}$; then $C_{\phi_a}(z^ng) = z^nG$ for some $G \in (\Theta H^2)^{\perp}$. This means that
$$(az + 1 - a)^n g \circ \phi_a(z) = z^n G(z) \,\ \,\ \,\ \forall z \in \mathbb{D}.$$
    Evaluating at $0$ we conclude that $g(1 - a) = 0$. But not every function in $(\Theta H^2)^{\perp}$ satisfies this condition: the reproducing kernel $\kappa_{0}^{\Theta}$ given by $\kappa_{0}^{\Theta}(z) = 1 - \overline{\Theta(0)} \Theta(z)$ is such that $\kappa_{0}^{\Theta}(1 - a) = 1 -  \overline{\Theta(0)} \Theta(1 - a) \neq 0$ (because $\Theta$ is inner and non constant, which implies that $\overline{\Theta(0)}$ and $ \Theta(1 - a)$ are in $\mathbb{D}$).
\end{proof}

\begin{cor}
   If $S$ is inner singular and $n \in \mathbb{N}$ then $(z^n S H^2)^{\perp}$ is the direct sum of an invariant and a non-invariant $C_{\phi_a}$-subspace.
\end{cor}
\begin{proof}
It is known that 
    $$(z^n S H^2)^{\perp} = (z^nH^2)^{\perp} \oplus z^n (SH^2)^{\perp}$$
    (see \cite[Lemma 14.6]{fricain1}). The result follows from the proposition above.
\end{proof}

To end this section, we will present some consequences involving universality. Since the operator $C_{\phi_a}$ has universal translates (see \cite{noor}), the minimal elements of $\mathrm{Lat}  (C_{\phi_a})$ are interesting from the point of view of the ISP.  

\begin{cor} If a model space is minimal and $C_{\phi_a}$-invariant, then it is one-dimensional.
\end{cor}
\begin{proof}
The model spaces $(e_nH^2)^{\perp}$ in Theorem \ref{mp} are minimal precisely when $n=1$, in which case they consist of only the constant functions.
\end{proof}

A more general question inspired by this corollary is what happens if some minimal  cyclic invariant subspace $K_f:=\overline{\mathrm{span}}(C_{\phi_a}^nf)_{n\in\mathbb{N}}$ contains a function $g \neq 0$ that belongs to some model space. In this case, of course $K_g = K_f$ due to minimality. We get a complete description of such $K_f$ in the following situation.

\begin{prop}
 Suppose that $\Theta$ is a non-constant inner function such that $1 \notin \sigma(\Theta)$ and let $f \in (\Theta H^2)^{\perp}$. Then $K_f$ is minimal if and only if $dim \,\ K_f = 1$.
\end{prop}

\begin{proof}
     Since $\sigma(\Theta)$ is a compact subset of $\mathbb{D}$ and $1 \notin \sigma(\Theta)$ we can take an arc $\mathbb{I}$ around $1$ such that $\mathbb{I} \cap \sigma(\Theta) = \emptyset$. It is known that each function $f \in (\Theta H^2)^{\perp}$ has an analytic continuation across $S^1 - \sigma(\Theta)$, in particular, each $f$ has an analytic continuation across  $\mathbb{I}$ and then is analytic at $1$ (see \cite{fricain1}, Lemma 14.27). The result follows from Corollary $4.4$ in \cite{noor}.
\end{proof}

\section{Invariant Beurling subspaces}

If $\Theta$ is a zero free inner function then $\Theta = \lambda S_{\mu}$ for some unimodular constant and $S_{\mu}$ a singular inner function. In this case, by Proposition \ref{spectrainner} we conclude that $\sigma(\Theta) = \mathrm{supp} (\mu) \subseteq S^1$. If $\mathrm{supp}(\mu)$ has only one point $\xi_0 \in S^1$ then
$$S_{\mu}(z) = e^ {- K \left(\frac{\xi_0 + z}{\xi_0 - z}\right)}$$
where $K = \mu(\{\xi_0\})$. Our goal in this section is to understand when a Beurling subspace is invariant under $C_{\phi_a}$. The following characterization of $C_\phi$-invariant Beurling subspaces induced by Blaschke products was provided by Cowen and Wahl for elliptic non-automorphisms \cite{Cowen-Wahl} and more generally by Bose, Muthukumar and Sarkar \cite{Beurlingtype}.
\begin{thm}\label{Blashke invariance} Let $B$ be a Blaschke product and let $\phi$ be a holomorphic self-map of $\mathbb{D}$. Then the following statements are equivalent:
\begin{enumerate}
    \item $BH^2$ is $C_{\phi}$-invariant.
    \item $mult_{B}(w) \leq mult_{B \circ \phi}(w)$ for every $w \in Z(B)$.
\end{enumerate}
\end{thm}

If   \[
B(z) =  z^N \prod\limits_{i = 1}^{\infty} \frac{\overline{z_i}}{|z_i|} \left( \frac{z_i - z}{ 1 - \overline{z_i}z } \right) \,\ \,\ 
\]
is any Blaschke product, then
\[
B \circ \phi_a(z) =  (az + 1 - a)^N \prod\limits_{i = 1}^{\infty} \frac{\overline{z_i}}{|z_i|} \left( \frac{z_i - az - 1 + a}{ 1 - \overline{z_i}(az + 1 - a) } \right) .
\]

Consider the example of a Blaschke product formed by the zeros $(1 - a^{2n})_{n \in \mathbb{N}}$. Every $w \in Z(B)$ has multiplicity equal to $1$ and $B \circ \phi_{a^2}(1 - a^{2n}) = B(1 - a^{4n}) = 0$. This means that the zeros of $B$ are all zeros of $B \circ \phi_{a^2}$ which implies that $BH^2$ is $C_{\phi_{a^2}}$-invariant by Theorem \ref{Blashke invariance}. On the other hand $B \circ \phi_a(1 - a^2) = B(1 - a^3) \neq 0$ and thus using Theorem \ref{Blashke invariance} again we conclude that $B H^2$ is not $C_{\phi_a}$-invariant. In particular, this example demonstrates that not all invariant Beurling subspaces for some $C_{\phi_a}$ are necesarily invariant under the entire family $\{C_{\phi_a}:a\in(0,1)\}$. We state this for use later.



   

\begin{exem}\label{common}
    The Blaschke product $B$ formed by the sequence of zeros $(1 - a^{2n})_{n \in \mathbb{N}}$ all with multiplicity $1$ is such that $BH^2$ is $C_{\phi_{a^2}}$-invariant, but is not $C_{\phi_{a}}$-invariant.
\end{exem}

In general, if $z_0 \in \mathbb{D}$ then the Blaschke product $B$ formed by the simple zeros $z_0, az_0 + 1 - a, a^2z_0 + 1 - a^2 \ldots$ is such that $BH^2$ is $C_{\phi_a}$-invariant.  This lead us to the following conclusion:
\begin{cor}\label{bminimal}
    Let $\Theta$ be an inner function such that $\Theta H^2$ is $C_{\phi_a}$-invariant. Then there exists an inner function $\Upsilon $ with $\Upsilon H^2 \subsetneq \Theta H^2$ and $\Upsilon H^2\in\mathrm{Lat}(C_{\phi_a})$. In particular, no minimal invariant subspace for $C_{\phi_a}$ can be a Beurling subspace.
\end{cor}
\begin{proof}
    Let $z_0 \in \mathbb{D}$ and consider the Blaschke product $B$ formed by the sequence $ (a^nz_0 + 1 - a^n)_{n\geq 0}$. By the above discussion $BH^2$ is $C_{\phi_a}$-invariant. If we let $\Upsilon = B\Theta$, then $\Upsilon H^2 \subsetneq  \Theta H^2$. To prove that $\Upsilon H^2$ is invariant under $C_{\phi_a}$, we argue as follows. By \cite[Theorem 2.3]{Beurlingtype} we have  $\frac{B \circ \phi_a}{B} \in \mathcal{S} \subseteq H^{\infty}$. Moreover, if $\Upsilon g = B\Theta g \in \Upsilon H^2$ then
$$\frac{(B \circ \phi_a )}{B} B(\Theta \circ \phi_a) (g \circ \phi_a) = \frac{(B \circ \phi_a )}{B} B \Theta g_2 =   B  \Theta \frac{(B \circ \phi_a )}{B} g_2 = \Upsilon h.$$
    where $h \in H^2$. Thus $C_{\phi_a}(\Upsilon g) = \Upsilon h \in \Upsilon H^2$ and we are done. The last claim follows directly by the definition of minimality.
   \end{proof}
Our main goal in this section is to study the $C_{\phi_a}$-invariant Beurling subspaces. We start with the following lemma whose proof is partly inspired by ideas contained within \cite[Section 7]{Cesaro 2}.

\begin{lem}\label{mainbeurling}
    Suppose $\Theta$ is a nonconstant inner function and $\Theta H^2$ is $C_{\phi_a}$-invariant for some $a\in(0,1)$. Then $1 \in \sigma(\Theta)$. If $1\notin \overline{Z(\Theta)}$, then $\sigma(\Theta) \cap S^1=\{1\}$ and $1 \in supp(\mu)$, where $\mu$ is the singular measure associated with $\Theta$.
       
\end{lem}
\begin{proof} Suppose that $1 \notin \sigma(\Theta)$. By (\cite{Beurlingtype}, Theorem $2.3$) $\Theta H^2$ is $C_{\phi_a}$-invariant if and only if
\begin{equation}\label{star}
        \frac{\Theta \circ \phi_a}{\Theta} \in \mathcal{S}(\mathbb{D})
          \end{equation}
where $\mathcal{S}(\mathbb{D}) = \{f \in H^{\infty} \,\ | \,\ \|f\|_{\infty} \leq 1\}.$  Since $1 \notin \sigma(\Theta)$ we know that $\Theta$ has an analytic continuation across some open arc $I$ in $S^1$, $1 \in I$ and $|\Theta(w)| = 1$ for all $w \in I$. Now, consider the sequence $(1 - a^n)_{n \in \mathbb{N}} = \{\phi_{a^n}(0)\}_{n \in \mathbb{N}}$. Note that there exists $n_0 \in \mathbb{N}$ such that for every $n \geq n_0$, we have $|\Theta(1 - a^n)| \geq \frac{1}{2}$, otherwise we conclude that for some subsequence $(1 - a^{n_k})$ we have $\frac{1}{2} \geq |\Theta(1 - a^{n_k})| \to |\Theta(1)| = 1$ by continuity and this is a contradiction. So, for $n \geq n_0$ the condition \eqref{star} implies
\[
\ldots \leq  |\Theta (1 - a^{n_{0} + 2})| \leq |\Theta (1 - a^{n_{0} + 1})| \leq |\Theta(1 - a^{n_0})| \leq 1 
\]
when we evaluate $\frac{\Theta \circ \phi_a}{\Theta}$ at the points $(1 - a^{n})_{n \geq n_0}$. Thus
\begin{equation}\label{diamond}
    |\Theta (1 - a^{n_{0}+ j })| \leq |\Theta(1 - a^{n_0})| \leq 1. 
\end{equation} 
Since $(1 - a^{n_0 + j}) \to 1$, letting $j \to \infty$ in \eqref{diamond} gives $|\Theta(1 - a^{n_0})| = 1$. This implies $\Theta$ must be constant since $\Theta(\mathbb{D})\subset\mathbb{D}$ for non-constant inner functions. This contradiction  proves that $1 \in \sigma(\Theta)$. To prove the second part, we start by writing $\sigma(\Theta) = \overline{\{z_1, z_2, \ldots\,\}} \cup \mathrm{supp}(\mu)$ where $(z_n)_{n \in \mathbb{N}}$ are the zeros of $\Theta$. Let $\xi \in S^1 - \{1\}$ such that $\xi \in \sigma(\Theta)$. By Proposition \ref{spectrainner} we obtain
\begin{align*}
   \liminf\limits_{w \to \xi, \,\  w \in \mathbb{D}} |\Theta(w)| = 0 .
\end{align*}
 As $1\notin \overline{Z(\Theta)}$, we can choose $m_0 \in \mathbb{N}$ such that $\Theta$ is zero free in $\phi_{a^{n}}(\mathbb{D})$ for all $n \geq m_0$. Since $\Theta H^2$ is $C_{\phi_a}$-invariant, it is also $C_{\phi_{a^{m_0+ 1}}}$-invariant.  By \cite[Theorem 2.3]{Beurlingtype} we can write $\Theta \circ \phi_{a^{m_0 + 1}} = \Theta g$ for some $g \in \mathcal{S}(\mathbb{D})$. Note that $a^{m_0 + 1}\xi + 1 - a^{m_0 + 1} \in \phi_{a^{m_0}}(\mathbb{D})$ (because $\xi \neq 1$) and 
 \begin{align*}
   \liminf\limits_{w \to \xi, \,\  w \in \mathbb{D}} |\Theta \circ \phi_a^{m_0 + 1}(w)| & =  \liminf\limits_{w \to \xi, \,\  w \in \mathbb{D}} |\Theta(a^{m_0 + 1}w + 1 - a^{m_0 + 1})| \\
   & = |\Theta(a^{m_0 + 1}\xi + 1 - a^{m_0 + 1})| \neq 0
\end{align*}
because $\Theta$ is zero-free in  $\phi_{a^{m_0}}(\mathbb{D})$. So
\begin{align*}
   \liminf\limits_{w \to \xi, \,\  w \in \mathbb{D}} |g(w)| = \infty
\end{align*}
 which implies that $g$ is unbounded and we arrive at a contradiction. So the only possible point in $\sigma(\Theta) \cap S^{1}$ is $1$. For the last claim, since $1 \notin  \overline{ \{z_1, z_2, \ldots \} }$ the only possibility is $1 \in \mathrm{supp}(\mu)$.
\end{proof}

It is known that the space $e^{-K ( \frac{1 + z}{1 -z} )}H^2$ (where $K > 0$ is a constant) is  $C_{\phi_a}$-invariant (see \cite[Theorem 6]{Cowen-Wahl}). As a consequence of this and Proposition \ref{mainbeurling}, we obtain the main result of this section.

\begin{thm}\label{dic}
    Let $a \in (0,1)$ and $\Theta \in H^2$ inner. If $\Theta H^2$ is $C_{\phi_a}$-invariant, then exactly one of the following occurs:
    \begin{itemize}
        \item $\Theta(z) = \lambda e^{-K ( \frac{1 + z}{1 -z} )}$ for some $K > 0$ and $|\lambda| = 1$, or
        \item $\Theta$ has infinitely many zeros accumulating at $1$.
    \end{itemize}
\end{thm}
\begin{proof}
    If $\Theta$ is zero free then $\sigma(\Theta) = \mathrm{supp}(\mu) = \{1\}$ by Lemma \ref{mainbeurling}, where $\mu$ is the measure associated to the singular inner part of $\Theta$. Thus $\Theta$ is as desired. If $\Theta$ has a zero $z_0$, as $\Theta H^2$ is $C_{\phi_a}$-invariant we have
     $$\frac{\Theta \circ \phi_a}{\Theta} \in \mathcal{S}(\mathbb{D}) \,\ \,\  
     $$
     (\cite{Beurlingtype}, Theorem $2.3$). So $\Theta(az_0 + 1 -a) = 0$ otherwise this quotient is not analytic. But $\Theta H^2$ is also $C_{\phi_{a^2}}$-invariant because $C_{\phi_{a^2}}(\Theta H^2) = C_{\phi_a}(C_{\phi_a}(\Theta H^2))\subseteq \Theta H^2$. Using the same argument again we conclude that
     $$\frac{\Theta \circ \phi_{a^2}}{\Theta} \in \mathcal{S}(\mathbb{D}) \,\ \,\  
     $$
     and thus $\Theta(a^2z_0 + 1 - a^2) = 0$. Repeating this argument for each $n \in \mathbb{N}$ we obtain $\Theta(a^nz_0 + 1 - a^n) = 0$ which implies the desired result.
\end{proof}

 Due to Example \ref{common} there exist Beurling spaces that are invariant under $C_{\phi_{a}}$ for some, but not all $a \in (0,1)$. Hence the second case above does occur. As a consequence we obtain a recent result of Gallardo-Gutiérrez, Partington and Ross \cite[Theorem 7.7]{Cesaro 2}.

  \begin{cor}\label{Cesaro Model}
      A model space $(\Theta H^2)^\perp$ is invariant under the Cesàro operator $\mathcal{C}$ if and only if $\Theta(z) = \lambda e^{-K ( \frac{1 + z}{1 -z} )}$ for some $K > 0$ and $|\lambda| = 1$.
  \end{cor}
\begin{proof} A model space $(\Theta H^2)^\perp$ is $\mathcal{C}$-invariant if and only if $\Theta H^2$ is $C_{\phi_a}$-invariant for all $a\in(0,1)$ by Theorem \ref{GP1}. By the proof of Theorem \ref{dic}, if $\Theta$ vanishes at some $z_0\in\mathbb{D}$ then $\Theta(az_0 + 1 - a) = 0$ for all $a \in (0,1)$ which is impossible unless $\Theta\equiv 0$. Hence $\Theta$ must be an atomic singular inner function.
\end{proof}

\section*{Acknowledgements}
This work constitutes a part of the doctoral thesis of the first author, partially supported by the Conselho Nacional de Desenvolvimento Cient\'{i}fico e Tecnol\'{o}gico - CNPq Brasil, under the supervision of the second named author.

\bibliographystyle{amsplain}

\end{document}